OpenStudy (anonymous):

find if the seq converges {(1+cosn)/n^2}

OpenStudy (tkhunny):

Hint: \(-1\le\cos(n)\le 1\)

OpenStudy (anonymous):

use the squeeze theorem?

OpenStudy (tkhunny):

Let's see you do that.

OpenStudy (anonymous):

\[0\le1+\cos n \le2\]

OpenStudy (tkhunny):

That's it. Now, what do you conclude?

OpenStudy (anonymous):

\[\le\frac{ 0 }{ n^2 }\le \frac{ 1+\cos n }{ n^2}\le \frac{ 2 }{ n^2 }\]

OpenStudy (anonymous):

converges to zero

OpenStudy (tkhunny):

Done! Good work.

OpenStudy (anonymous):

awesome thx

OpenStudy (abb0t):

That's comparison theory, not squeeze.

OpenStudy (zarkon):

No..it is the squeeze theorem